# Minimal codes from hypersurfaces in even characteristic

### v3

@@ 4 The cutting gap
In this section we consider the extension of the notion of being a cutting set to subspaces of codimension higher than $1$ given by the notion of $k$-fold strong blocking set, introduced first in [11]. This is also related to the geometry of higher Hamming weights, besides the applications considered in the aforementioned paper.
In the following, we will use projective dimensions and put $\dim(\emptyset)=-1$.
@@
Let $\Omega$ be a non-empty set of points in $\mathrm{PG}(r,\mathbb{K})$. Let $0\leq k\leq r$, we define the $k$-th cutting gap of $\Omega$ as the number
$\tau_{k}(\Omega)=k-\min\{\dim(\langle\Pi\cap\Omega\rangle)\colon\dim(\Pi)=k\}.$
According to the definitions of [11], $\Omega$ is a $k$-fold strong blocking set if and only if $\tau_{k}(\Omega)=0$. In this paper we shall call this property also as being $k$-cutting and consequently say that $\Omega$ is a $k$-cutting set if $\tau_{k}(\Omega)=0$. For some constructions of sets which are $k$-cutting for some chosen values of $k$ we refer to [11, §3].
It is immediate to see that $\Omega$ is a cutting set or, more in detail, a cutting set with respect to the hyperplanes when $\tau_{r-1}(\Omega)=0$. In particular, $\Omega$ is $k$-cutting if any $k$-dimensional space is spanned by intersection with $\Omega$.
@@
Let $\Omega$ be a $k$-cutting set of $\mathrm{PG}(r,\mathbb{K})$. Then $\Omega$ is also $\ell$-cutting for all $k\leq\ell\leq r$.
@@
If $k=r$ there is nothing to prove. Otherwise it is enough to prove that $\tau_{k}(\Omega)=0$ implies $\tau_{k+1}(\Omega)=0$. Any subspace $\Pi$ of dimension $k+1$ can be regarded as the union of all of the subspaces of dimension $k$ contained in it. So,
@@
The thesis follows. ∎
We can easily write all the cutting gaps of non-degenerate Hermitian varieties using the properties of the induced polarity.
@@
$\tau_{r-t}(\mathscr{H}^{r})=\begin{cases}0&\textnormal{ for }t>\lfloor\frac{r}{2}\rfloor,\\ 1&\textnormal{ for }t\leq\lfloor\frac{r}{2}\rfloor.\end{cases}$
@@
Since Hermitian varieties over finite fields defined over a projective space always have at least one rational point, it is immediate to see that any Hermitian variety in a subspace $\Pi=\mathrm{PG}(V_{k+1})$ of dimension $k$ in $\mathrm{PG}(r,q^{2})$ is a blocking set for the lines of the space; so it is either a hyperplane of $\Pi$ or it spans $\Pi$. In the former case, the matrix of the Hermitian form restricted to $V_{k+1}\times V_{k+1}$ has rank $1$; in the latter it has either rank 00 (and it contains $\Pi$), or it has rank at least $2$.
Denote now by $H_{r+1}$ the non-singular Hermitian matrix describing the variety $\mathscr{H}^{r}$. Observe that for any subspace $\Pi$ of $\mathrm{PG}(r,q^{2})$ of codimension $t$, the form induced by $H_{r+1}$ on $\Pi$ has rank at least $r+1-2t$, since it corresponds to a $(r+1-t)\times(r+1-t)$ minor of $H_{r+1}$. So, if $r+1-2t>1$, that is $t<r/2$, this form is guaranteed to have rank at least $2$; thus, by the first part of the proof, $\Pi$ meets $\mathscr{H}^{r}$ in a (possibly degenerate) Hermitian variety spanning it. It follows that $\mathscr{H}^{r}$ is $(r-t)$-cutting and $\tau_{r-t}(\mathscr{H}^{r})=0$. On the other hand, if $t\geq r/2$ there exist spaces $\Pi$ meeting $\mathscr{H}^{r}$ in just a hyperplane of $\Pi$; this shows that $\mathscr{H}^{r}$ is not $(r-t)$-cutting for $t\leq r/2$ and $\tau_{r-t}(\mathscr{H}^{r})=1$. ∎
We leave as an open problem for a future work to determine if the list of cutting gaps of Theorem 4.3 is a characterization of Hermitian varieties among the family of quasi-Hermitian ones.
Being a $2$-fold blocking set for the lines implies automatically to be a $2$-fold blocking set. In particular, in $\mathrm{PG}(2,q)$ the only possible value of $k$ is $1$ and $\tau_{1}(\Omega)=0$ if and only if $\Omega$ is a $2$-fold blocking set for the lines, while $\tau_{1}(\Omega)=1$ if $\Omega$ is a blocking set and it is not a $2$-fold blocking set. Finally, $\tau_{1}(\Omega)=2$ if there are external lines to $\Omega$. This determines provides lower bounds on the possible sizes of $\Omega$.
A trivial observation about the cutting gaps is that $\tau_{k}(\Omega)>k$ if and only if there are $k$-subspaces disjoint from $\Omega$. In this case $\tau_{k}(\Omega)$ does not provide any further insight on the geometry of the point-set $\Omega$. In order to get a more refined view, we here also propose a definition of modified cutting gaps. Suppose $-1\leq s<k$, we define
$\tau_{k,s}(\Omega)=k-\min\{\dim(\langle\Pi\cap\Omega\rangle)\colon\dim(\Pi)=k,\dim(\langle\Pi\cap\Omega\rangle)\geq s\}.$
By definition $\tau_{k,-1}(\Omega)=\tau_{k}(\Omega)$ is the same as the cutting gap introduced before. Also $\tau_{k,s}(\Omega)\leq\tau_{k,s^{\prime}}(\Omega)$ whenever $s>s^{\prime}$, so the list of values of $\tau_{k,s}(\Omega)$ is non-increasing in $s$ (for a fixed $k$). It is also obvious whenever the point set $\Omega$ spans at least an $s$-space (but we usually assume $\Omega$ to span at least a $k$-space, otherwise some values are trivial), $\tau_{k,s}(\Omega)\leq k-s$. The expected general behavior is that equality holds, i.e. $\tau_{k,s}(\Omega)=k-s$, but, as the following example shows, this is not always the case.
@@
Let $\Omega:=\mathscr{Q}^{-}(3,q)\setminus\{P\}$ be the set of all points of an elliptic quadric with the exclusion of a point $P$ in $\mathrm{PG}(3,q)$. Then $\tau_{2}(\Omega)=\tau_{2,-1}(\Omega)=3$, as the tangent plane to $\mathscr{Q}^{-}(3,q)$ at $P$ is disjoint from $\Omega$. The tangent plane to $\mathscr{Q}^{-}(3,q)$ at any point $Q$ different from $P$ meets $\Omega$ just in $Q$, so $\tau_{2,0}(\Omega)=2$. Finally, observe that any plane meeting $\Omega$ in at least two distinct points is necessarily non-tangent to $\mathscr{Q}^{-}(3,q)$, so it meets $\Omega$ either in a conic or in a conic minus the point $P$; hence $\tau_{2,1}(\Omega)=0$.
We leave also to future works to investigate in more detail how to characterize the properties of the modified cutting gaps as well as to describe sets where $\tau_{k,s}<k-s$ for at least some values of $k$ and $s$.
In the following section, we will study the (modified) cutting gap for the variety $\mathscr{V}^{r}_{\varepsilon}$ for $r=4$.
@@ 4.2 Intersection of $\mathscr{V}^{4}_{\varepsilon}$ with planes
The number of solutions of Equation (32) corresponds to the number of absolute points in $\mathrm{PG}(1,2^{2e})$ of a (possibly degenerate) $s$-sesquilinear form. By [12, Lemma 6] this corresponds to either $0,1,2$ or $t$ points where $t$ is the cardinality of a subline over the fixed subfield of $s$. In particular, for $e\equiv 1\pmod{4}$ we have $t=5$ and for $e\equiv 3\pmod{4}$ we have $t=3$. ∎

### v4

@@ 4 The cutting gap
Minimal projective linear codes and cutting blocking sets with respect to the hyperplanes are essentially the same kind of object, see [7] and [22]. In [13], the notion of $t$-fold strong blocking set $\Lambda$ is introduced, as a set of points in a projective space $\mathrm{PG}(n,q)$ such that every (t1)(t-1)-dimensional projective subspace is spanned by $t$ points in $\Lambda$. Clearly, being a cutting blocking set with respect to the hyperplanes and being a $n$-fold strong blocking set mean the same.
In this section we introduce the concept of cutting gap, as a measure of the amount in which a set $\Lambda$ fails to be a $t$-fold strong blocking set for some $t$. It provides a quantitative measure of how effectively a point set intersects $k$-dimensional subspaces in a projective space. This notion is closely related to the geometry of higher Hamming weights and has applications in coding theory and incidence geometry, as first explored in [13]. In particular, cutting gaps allow us to distinguish between point sets that merely intersect subspaces and those that span them, offering a refined tool for analyzing the structure of algebraic varieties.
Throughout this section, we use projective dimensions and adopt the convention $\dim(\emptyset)=-1$.
@@
Let $\Omega$ be a non-empty set of points in $\mathrm{PG}(r,\mathbb{K})$, and let $0\leq k\leq r$. The $k$-th cutting gap of $\Omega$ is defined as
$\tau_{k}(\Omega)=k-\min\left\{\dim\left(\langle\Pi\cap\Omega\rangle\right)\colon\dim(\Pi)=k\right\}.$
According to [13], $\Omega$ is a $(k+1)$-fold strong blocking set if and only if $\tau_{k}(\Omega)=0$. In this paper, we refer to this property as being $k$-cutting, and we say that $\Omega$ is a $k$-cutting set if $\tau_{k}(\Omega)=0$. For explicit constructions of $k$-cutting sets for selected values of $k$, see [13, §3].
Note that $\Omega$ is a classical cutting set with respect to hyperplanes precisely when $\tau_{r-1}(\Omega)=0$. More generally, $\Omega$ is $k$-cutting if every $k$-dimensional subspace is spanned by its intersection with $\Omega$.
@@
Let $\Omega$ be a $k$-cutting set in $\mathrm{PG}(r,\mathbb{K})$. Then $\Omega$ is also $\ell$-cutting for all $k\leq\ell\leq r$.
@@
If $k=r$, the statement is trivial. Otherwise, suppose $\tau_{k}(\Omega)=0$. Any $(k+1)$-dimensional subspace $\Pi$ can be expressed as the union of its $k$-dimensional subspaces:
@@
Hence, $\tau_{k+1}(\Omega)=0$, and the result follows by induction. ∎
We now compute the cutting gaps for non-degenerate Hermitian varieties using properties of the induced polarity.
@@
$\tau_{r-t}(\mathscr{H}^{r})=\begin{cases}0&\text{if }t>\left\lfloor\frac{r}{2}\right\rfloor,\\ 1&\text{if }t\leq\left\lfloor\frac{r}{2}\right\rfloor.\end{cases}$
@@
Hermitian varieties over finite fields always contain at least one rational point. Thus, any subspace $\Pi=\mathrm{PG}(V_{k+1})$ of dimension $k$ in $\mathrm{PG}(r,q^{2})$ intersects $\mathscr{H}^{r}$ in a blocking set for the lines of $\Pi$, which either spans $\Pi$ or is a hyperplane of it.
Let $H_{r+1}$ be the non-singular Hermitian matrix defining $\mathscr{H}^{r}$. For any subspace $\Pi$ of codimension $t$, the induced Hermitian form has rank at least r+12tr+1-2t, as it corresponds to a $(r+1-t)\times(r+1-t)$ minor of $H_{r+1}$. If r+12t>1r+1-2t>1, i.e., $t<r/2$, the form has rank at least $2$, and $\Pi\cap\mathscr{H}^{r}$ spans $\Pi$, so $\tau_{r-t}(\mathscr{H}^{r})=0$.
Conversely, if $t\geq r/2$, there exist subspaces $\Pi$ such that $\Pi\cap\mathscr{H}^{r}$ is a hyperplane of $\Pi$, implying $\tau_{r-t}(\mathscr{H}^{r})=1$. ∎
We leave as an open problem whether the list of cutting gaps in Theorem 4.3 characterizes Hermitian varieties among quasi-Hermitian ones.
In $\mathrm{PG}(2,q)$, the only relevant value is $k=1$. Then: $\tau_{1}(\Omega)=0$ if and only if $\Omega$ is a $2$-fold blocking set for the lines; $\tau_{1}(\Omega)=1$ if $\Omega$ is a blocking set but not $2$-fold; $\tau_{1}(\Omega)=2$ if there exist lines disjoint from $\Omega$. This provides lower bounds on the size of $\Omega$.
A trivial observation is that $\tau_{k}(\Omega)>k$ if and only if there exist $k$-subspaces disjoint from $\Omega$. In such cases, $\tau_{k}(\Omega)$ does not offer any further geometric insight. To refine this, we introduce the notion of modified cutting gaps.
Let $1\leq s<k$. Define:
$\tau_{k,s}(\Omega)=k-\min\left\{\dim\left(\langle\Pi\cap\Omega\rangle\right)\colon\dim(\Pi)=k,\dim(\langle\Pi\cap\Omega\rangle)\geq s\right\}.$
By definition, $\tau_{k,-1}(\Omega)=\tau_{k}(\Omega)$. Moreover, $\tau_{k,s}(\Omega)\leq\tau_{k,s^{\prime}}(\Omega)$ whenever $s>s^{\prime}$, so the sequence is non-increasing in $s$ for fixed $k$. If $\Omega$ spans at least an $s$-space, then $\tau_{k,s}(\Omega)\leq k-s$. While equality is expected in general, the following example shows that this is not always the case.
@@
Let $\Omega:=\mathscr{Q}^{-}(3,q)\setminus\{P\}$ be the set of points of an elliptic quadric in $\mathrm{PG}(3,q)$ with one point $P$ removed. Then: $\tau_{2}(\Omega)=\tau_{2,-1}(\Omega)=3$, since the tangent plane at $P$ is disjoint from $\Omega$; $\tau_{2,0}(\Omega)=2$, as tangent planes at other points intersect $\Omega$ in a single point; $\tau_{2,1}(\Omega)=0$, since any plane intersecting $\Omega$ in at least two points meets it in a conic or a conic minus $P$.
We leave to future work a deeper investigation of the properties of modified cutting gaps and the classification of point sets for which $\tau_{k,s}(\Omega)<k-s$ for some values of $k$ and $s$.
@@ 4.2 Intersection of $\mathscr{V}^{4}_{\varepsilon}$ with planes
The number of solutions of Equation (32) corresponds to the number of absolute points in $\mathrm{PG}(1,2^{2e})$ of a (possibly degenerate) $s$-sesquilinear form. By [14, Lemma 6] this corresponds to either $0,1,2$ or $t$ points where $t$ is the cardinality of a subline over the fixed subfield of $s$. In particular, for $e\equiv 1\pmod{4}$ we have $t=5$ and for $e\equiv 3\pmod{4}$ we have $t=3$. ∎